OpenStudy (anonymous):

Please! Could anyone please help me describe the sketch of the following curve, and explain why it looks this way?

OpenStudy (anonymous):

\[h:[0,1]\rightarrow C,where: h(t)=3\cos(2 \pi t) + 2i \sin(2 \pi t)\]